OpenStudy (anonymous):

tan(2 arc cot(1/2) )=

OpenStudy (anonymous):

Let arccot 1/2 be x then Tan 2x = 2 Tan x / 1-Tan^2 x and cot x = 1/2 -> tan x = 2 so 2*2/1-2^2 = -4/3